# Markov-type state models to describe non-Markovian dynamics

### v1

@@ Abstract
When clustering molecular dynamics (MD) trajectories into a few metastable conformational states, the Markov state models (MSMs) assumption of timescale separation between fast intrastate fluctuations and rarely occurring interstate transitions is often not valid. Hence, the naive estimation of the macrostate transition matrix via simply counting transitions between the states leads to significantly too short implied timescales and thus to too fast population decays. In this work, we discuss advanced approaches to estimate the transition matrix. Assuming that Markovianity is at least given at the microstate level, we consider the Laplace-transform based method by Hummer and Szabo, as well as a direct microstate-to-macrostate projection, which by design yields correct macrostate population dynamics. Alternatively, we study the recently proposed quasi-MSM ansatz of Huang and coworkers to solve a generalized master equations, as well as a hybrid method that employs MD at short times and MSM at long times. Adopting a one-dimensional toy model and an all-atom folding trajectory of HP35, we discuss the virtues and shortcomings of the various approaches.
@@ 1 Introduction
Classical molecular dynamics (MD) simulations are a versatile tool for gaining insights into complex biomolecular processes.1 To facilitate the interpretation of the ever-increasing amount of data obtained from such simulations, coarse grained models such as Langevin equations2, 3, 4 and Markov state models 5, 6, 7, 8 (MSMs) can be useful. Interpreting MD trajectories in terms of memoryless transitions between metastable conformational states, MSMs are popular because they provide a generally accepted state-of-the-art analysis, promise to predict long-time dynamics from short trajectories, and are straightforward to build using open-source packages. 9, 10, 11 The usual workflow to construct an MSM consists of (i) selection of suitable input coordinates or features, (ii) dimensionality reduction from the high-dimensional feature space to some low-dimensional space of collective variables, (iii) geometrical clustering of these low-dimensional data into microstates, (iv) dynamical clustering of the microstates into metastable conformational states (or macrostates), and (v) estimation of the transition matrix associated with these states. Here, we are concerned with the last step of the procedure, that is, we want to discuss various approaches (and propose new ways) to estimate the macrostate transition matrix ${\boldsymbol{T}}(t)$, containing the probabilities $T_{IJ}$ that the system jumps from state $J$ to state $I$ within some pre-chosen lag time. To this end we assume that steps (i) - (iv) resulted in a partitioning of $N$ macrostates, such that the MD data can be converted to a state trajectory $I(t)$. Most straightforwardly, we may obtain ${\boldsymbol{T}}(t)$ by simply counting the transitions from $J$ to $I$ within $t$. Denoting the time-dependent state vector by ${\boldsymbol{P}}(t)=(P_{1},\ldots,P_{N})^{T}$ with state probabilities $P_{I}$, the time evolution of the state model can then be written as
@@
While this expression does not involve a dynamical approximation (such as the Markov approximation discussed below), its practical use is limited by the fact that the calculation of ${\boldsymbol{T}}(t)$ is restricted to times $t\lesssim t_{\rm max}$, where $t_{\rm max}$ is the length of the MD trajectories. That is, Eq. (1) is in essence a state representation of the dynamics, but it does not allow any prediction beyond the timescale of the MD simulations.
@@
${\boldsymbol{T}}(t\!=\!m\tau)\approx{\boldsymbol{T}}^{m}(\tau),$ (2)
which only requires short MD trajectories (in principle as short as $\tau$) to estimate the transition matrix. Since the time evolution of the MSM is given in steps of $\tau$, the lag time defines the time resolution of the model, and therefore needs to be chosen shorter than the fastest dynamics of interest. This condition, however, is often in conflict with the above requirement of a long enough lag time to achieve Markovianity of the system, i.e., a constant transition matrix. Therefore, in practice one often ends up with a state partitioning that is only approximately Markovian at the desired time resolution, which raises the question of the optimal MSM that best approximates the dynamics in this case.
There are various ways to compute the MSM transition matrix ${\boldsymbol{T}}(\tau)$, in order to optimize the approximation of the true transition matrix ${\boldsymbol{T}}(t)$ in Eq. (2). For one, we may go beyond the simple counting scheme mentioned above and invoke our knowledge of the microstates the macrostates are build of.12, 13, 14, 15 This approach exploits that microstates are typically structurally homogeneous, and therefore require only a relatively short lag time to randomize and show Markovian behavior. Macrostates, on the other hand, typically contain numerous structurally different microstates that may be separated by free energy barriers, resulting in long lag times to reach Markovianity. Along these lines, Hummer and Szabo 13 proposed a valuable formulation to optimally project the microstate dynamics onto the macrostate dynamics, which was shown to achieve significantly improved Markovianity.
@@
where ${\boldsymbol{K}}(t)$ is the memory kernel matrix that accounts for the non-Markovian dynamics of the system. To obtain an equation of motion for the transition matrix, we may replace ${\boldsymbol{P}}(t)$ by ${\boldsymbol{T}}(t)$ using Eq. (1). While the estimation of the memory kernel for (typically noisy and under-sampled) MD data represents a non-trivial task,21, 22, 23 Huang and coworkers17 recently proposed a promising method termed quasi-MSM (qMSM), which was successfully applied to various systems.24, 25
In this work, we discuss the theoretical basis and the performance of various approaches to estimate ${\boldsymbol{T}}(t)$, including the Hummer-Szabo projection,13 the qMSM ansatz of Huang and coworkers,17 and two new approaches termed ‘microstate-based’ and ‘hybrid MD/MSM’ method, respectively. Adopting a one-dimensional toy model and an all-atom folding trajectory of HP35,26 we discuss the applicability as well as the virtues and shortcomings of these approaches.
2 Theory
@@
2.1 From microstate to macrostate dynamics
In a typical MSM workflow we first construct (say, $n\!\sim\!10^{2}$ $10^{3}$) microstates from the MD data, which are subsequently lumped into a few (say, $N\lesssim 10$) macrostates. As explained in the Introduction, we assume that the microstates show Markovian behavior at the chosen lag time $\tau$, which is sufficiently short to resolve the fastest dynamics of interest. Hence, the microstate transition matrix ${\boldsymbol{t}}=\{t_{ij}\}$ satisfies the Chapman-Kolmogorov relation
@@ 2.1 From microstate to macrostate dynamics
This microstate-based evaluation of the macrostate transition matrix preserves all dynamical properties established for the microstates. For example, it yields by design exact macrostate population dynamics, and also provides correct state-to-state waiting times and transition pathways from a microstate Monte Carlo Markov chain simulations with subsequent projection on the macrostates. Compared to the local-equilibrium approximation that yields a constant transition matrix ${\boldsymbol{T}}_{\rm LE}(\tau)$ and the MSM relation (8), however, the microstate-based evaluation results in a time-dependent macrostate transition matrix.
@@
Inserting the resulting ${\boldsymbol{K}}_{n}$ in Eq. (12), we obtain the desired evolution of ${\boldsymbol{T}}(t)$ for long times ($t\gg\tau_{\mathrm{K}}$). Hence the qMSM formulation facilitates a straightforward approximate solution of an in principle exact generalized master equation, and also explicitly provides the memory matrix ${\boldsymbol{K}}(t)$ that accounts for the non-Markovian dynamics of the system. Requiring only MD data of length $t\approx\tau_{\mathrm{K}}$, qMSM allows us to predict long-time dynamics from short trajectories.
@@
Another issue is that qMSM keeps the additional term $\dot{{\boldsymbol{T}}}(0){\boldsymbol{T}}(t)$, which in principle vanishes for symmetry reasons,16 but may be non-zero in a numerical evaluation. Interestingly, this term exactly cancels the memory kernel at $t\!=\!0$ [which can be shown by calculating ${\boldsymbol{T}}_{n}$ and ${\boldsymbol{K}}_{n}$ for $n=0$ from Eqs. (12) and (13), respectively]. While this is again a consequence of the chosen discretization, it seems odd, because ${\boldsymbol{K}}_{0}$ is the only non-zero term in the Markov limit. In all cases considered, however, the neglect of the term seems to hardly change the numerical results. Another practical issue is that the calculation of the memory matrix from Eq. (13) can be seriously plagued by noise resulting from the MD data and the time derivatives. As a remedy, recently an integrative ansatz 25 and a time-convolutionless approach 20 of the generalized master equation were proposed.
@@ 2.3 Hybrid MD/MSM formulation
In the Introduction, we opposed in Eqs. (1) and (2) two standard ways to estimate the macrostate transition matrix ${\boldsymbol{T}}(t)=\{T_{IJ}(t)\}$. On the one hand, we may use the MD data to count the transitions from state $J$ to state $I$ within time $t$. The resulting transition matrix ${\boldsymbol{T}}_{\rm MD}(t)$ generates the correct time evolution of the state model according to Eq. (1), but is limited to times $t\approx t_{\rm max}$ with $t_{\rm max}$ being the length of the MD trajectories (and actually even some hundreds MD frames less to facilitate a sufficient time average). On the other hand, we may invoke the Markov approximation to obtain ${\boldsymbol{T}}(t\!=\!m\tau)\approx{\boldsymbol{T}}^{m}(\tau)$, which assumes that the transition matrix ${\boldsymbol{T}}(t)$ becomes constant for $t\gtrsim\tau$. While this only requires short MD trajectories, we need to choose a lag time $\tau$ that is long enough to achieve Markovianity, but at the same time short enough to resolve the fastest dynamics of interest.
@@
In this way, we have at short times the approximation-free transition matrix ${\boldsymbol{T}}_{\rm MD}(t)$, including full time resolution (as given by the MD data). At long times, we typically can allow for a rather long lag time.
@@ 3 Results
To compare the performance of the different methods, we test them on two model systems: A one-dimensional toy model lumping four microstates into two macrostates, and a recently established27 benchmark MSM of the folding of villin headpiece (aka HP35) using a long MD trajectory by Piana et al.26
@@ 3.1 One-dimensional toy model
${\boldsymbol{T}}_{\text{LE}}(t\!=\!n\tau_{0})=\begin{bmatrix}1-h/2&h/2\\ h/2&1-h/2\end{bmatrix}^{n}.$ (17)
@@
The Hummer-Szabo projection [Eq. (10)] gives ${\boldsymbol{T}}_{\rm HS}(t\!=\!n\tau_{0})={\boldsymbol{T}}_{\rm HS}^{n}$ with 13
@@
Figure 1: One-dimensional toy model. (a) Schematic free energy landscape, indicating four microstates 1, 2, 3 and 4, which are lumped into two macrostates $L$ and $R$. Depending on the ratio of the interstate transition probability h and the intrastate transition probability k, the macrostate dynamics is Markovian ($h/k\ll 1$, dark gray) or non-Markovian ($h/k\gtrsim 1$, dashed). (b) Implied timescale $t_{\rm I}$ of the two-macrostate model for $h/k=1$. Compared are the reference result obtained from the microstates (dashed black), the local equilibrium approximation (black), and the Hummer-Szabo projection (green), as well as the time evolution the result from qMSM (using the Kernel time $\tau_{\rm K}=3$, red) and of the microstate-based result (blue). (c) Normalized implied timescales obtained from the various methods, drawn as a function of the Markovianity parameter $h/k$. Time evolution of the macrostate population of the various methods, shown for (d) $h/k=1$ and (e) $h/k=5$. Also shown are results of the hybrid MD/MSM calculation [Eq. (14)] with $t_{\rm max}=10$. (f) Normalized memory kernel $K(t)$ of the qMSM method, obtained for $h/k=0.2$ (blue), $h/k=1$ (red), and $h/k=5$ (green).
As a first test, we wish to study how the various formulations reproduce the reference implied timescale $t_{\rm micro}$ of the microstates. Choosing $k=h=0.1$ (weakly non-Markovian case), Fig. 1b shows that the local equilibrium approximation $t_{\rm LE}$ underestimates $t_{\rm micro}$ almost by a factor 2, while the Hummer-Szabo result is only shorter by 13 %. The time evolution of ${\boldsymbol{T}}_{\text{qMSM}}(t)$ and ${\boldsymbol{T}}_{\text{Mic}}(t)$ reveals that the qMSM result converges rapidly to a value that is 29 % shorter as $t_{\rm micro}$, and that the result of the microstate-based method gradually converges to the correct result. We now choose $t=100$ (where qMSM and microstate-based results are virtually constant), and vary the Markovianity parameter $h/k$ from $h/k\ll 1$ (Markovian case) to $h/k\gtrsim 1$ (non-Markovian case). While for $h=0.1,\,k=0.5$ all methods yield the correct microstate timescale $t_{\rm micro}$ (except for $t_{\rm LE}$), the results vary considerably in the non-Markovian case. Choosing $h=0.5,\,k=0.1$, for example, $t_{\rm LE}$ underestimates $t_{\rm micro}$ by 87 %, $t_{\rm HS}$ by 40 %, and $t_{\rm qMSM}$ by 27 %, while the microstate-based result closely matches the reference.
@@
Choosing $h/k=1$ and $h/k=5$, Figs. 1d and 1e compare the population $P(t)$ obtained from the various formulations. While for $h/k=1$ all methods are quite close to the reference results, we find significant deviations in the strongly non-Markovian case $h/k=5$, where $P(t)$ in fact reveals two timescales. That is, the population exhibits a strong initial decay to $P(1)=0.75$ during the first time step, reflecting fast transitions over the shallow barrier at $x=0$ (Fig. 1a). All formulations reproduce this value correctly, except for the long-time approximation $P_{\rm HS}(t)$, which coincides with the reference for $t\gtrsim 7$. Being a short-time approximation, $P_{\rm LE}(t)$ decays significantly too fast for $t>1$. The qMSM calculation represents a clear improvement over these approximations and reproduces the exact results closely. Finally, we also show the hybrid calculation [Eq. (14)], which uses until $t_{\rm max}=10$ the exact transition matrix ${\boldsymbol{T}}_{\rm M}(t)$, and for longer times the local approximation ${\boldsymbol{T}}_{\rm LE}(t=nt_{\rm max})$. For this choice of $t_{\rm max}$, the hybrid result virtually matches the reference at all times.

### v2

@@ Abstract
When clustering molecular dynamics (MD) trajectories into a few metastable conformational states, the assumption of timescale separation between fast intrastate fluctuations and rarely occurring interstate transitions is often not valid. Hence, when we construct a Markov state model (MSM) from these states, the naive estimation of the macrostate transition matrix via simply counting transitions between the states may lead to significantly too short implied timescales and thus to too fast population decays. In this work, we discuss advanced approaches to estimate the transition matrix. Assuming that Markovianity is at least given at the microstate level, we consider the Laplace-transform based method by Hummer and Szabo, as well as a direct microstate-to-macrostate projection, which by design yields correct macrostate population dynamics. Alternatively, we study the recently proposed quasi-MSM ansatz of Huang and coworkers to solve a generalized master equation, as well as a hybrid method that employs MD at short times and MSM at long times. Adopting a one-dimensional toy model and an all-atom folding trajectory of HP35, we discuss the virtues and shortcomings of the various approaches.
@@ 1 Introduction
Classical molecular dynamics (MD) simulations are a versatile tool for gaining insights into complex biomolecular processes.1 To facilitate the interpretation of the ever-increasing amount of data obtained from such simulations, coarse grained models such as Langevin equations2, 3, 4 and Markov state models 5, 6, 7, 8 (MSMs) can be useful. Interpreting MD trajectories in terms of memoryless transitions between metastable conformational states, MSMs are popular because they provide a generally accepted state-of-the-art analysis, promise to predict long-time dynamics from short trajectories, and are straightforward to build using open-source packages. 9, 10, 11 The usual workflow to construct an MSM consists of (i) selection of suitable input coordinates or features, (ii) dimensionality reduction from the high-dimensional feature space to some low-dimensional space of collective variables, (iii) geometrical clustering of these low-dimensional data into microstates, (iv) dynamical clustering of the microstates into metastable conformational states (or macrostates), and (v) estimation of the transition matrix associated with these states.
Here, we are concerned with the last step of the procedure, that is, we want to discuss various approaches (and propose new ways) to estimate the macrostate transition matrix ${\boldsymbol{T}}(t)$. Since these formulations may go beyond the scope of a common MSM, we call them ’Markov-type state models.’ In what follows, we assume that steps (i) - (iv) resulted in a partitioning of $N$ macrostates, such that the MD data can be converted to a state trajectory $I(t)$. Moreover we assume that all states are connected such that the MSM is ergodic, and that the MD data represent a stationary and time-homogeneous process. To focus on the effects of non-Markovianity, we for now also assume sufficient sampling of the MD data (but see Sec. 3.2).
The macrostate transition matrix ${\boldsymbol{T}}(t)$ contains the probabilities $T_{IJ}$ that the system jumps from state $J$ to state $I$ within some pre-chosen lag time, which can be obtained by simply counting the transitions from $J$ to $I$ within $t$. When we denote the time-dependent state vector by ${\boldsymbol{P}}(t)=(P_{1},\ldots,P_{N})^{T}$ with state probabilities $P_{I}$, the time evolution of the state model can be written as
@@
We note that the time-dependent transition matrix ${\boldsymbol{T}}(t)$ does not involve a dynamical approximation (such as the Markov approximation discussed below), but is directly obtained from the MD data. As a consequence, its practical use is limited by the fact that the calculation of ${\boldsymbol{T}}(t)$ is restricted to times $t\lesssim t_{\rm max}$, where $t_{\rm max}$ is the length of the MD trajectories. That is, Eq. (1) is in essence a state representation of the dynamics, but it does not allow any prediction beyond the timescale of the MD simulations.
@@
${\boldsymbol{T}}(t\!=\!m\tau)\approx{\boldsymbol{T}}^{m}(\tau)$ (2)
[…]
There are various ways to compute the MSM transition matrix ${\boldsymbol{T}}(\tau)$, in order to optimize the approximation of the true transition matrix ${\boldsymbol{T}}(t)$ in Eq. (2). For one, we may go beyond the simple counting scheme mentioned above and invoke our knowledge of the microstates the macrostates are built of.12, 13, 14, 15 This approach exploits that microstates are typically structurally homogeneous, and therefore require only a relatively short lag time to randomize and show Markovian behavior. Macrostates, on the other hand, typically contain numerous structurally different microstates that may be separated by free energy barriers, resulting in long lag times to reach Markovianity. Along these lines, Hummer and Szabo 13 proposed a valuable formulation to optimally project the microstate dynamics onto the macrostate dynamics, which was shown to achieve significantly improved Markovianity.
@@
where ${\boldsymbol{K}}(t)$ is the memory kernel matrix that accounts for the non-Markovian dynamics of the system. To obtain an equation of motion for the transition matrix, we may replace ${\boldsymbol{P}}(t)$ by ${\boldsymbol{T}}(t)$ using Eq. (1). While the estimation of the memory kernel for (typically noisy and under-sampled) MD data represents a non-trivial task,21, 22, 23 Huang and coworkers17 recently proposed a promising method termed quasi-MSM (qMSM), which was successfully applied to various systems.24, 25 We also wish to mention a simpler but often effective way of taking memory into account. By ‘coring’ the macrostate trajectory, we either request that a transition from one state to another must reach some pre-defined core region of the other state,5, 26, 27 or that the system spends a certain minimum time in the new state 28. The latter approach, termed ‘dynamical coring’ was shown to considerably improve the Markovianity of the resulting metastable states.29
In this work, we discuss the theoretical basis and the performance of various approaches to estimate ${\boldsymbol{T}}(t)$, including the Hummer-Szabo projection,13 the qMSM ansatz of Huang and coworkers,17 and two new approaches termed ‘microstate-based’ and ‘hybrid MD/MSM’ method, respectively. Adopting a one-dimensional toy model and an all-atom folding trajectory of HP35,30 we discuss the applicability as well as the virtues and shortcomings of these approaches.
2 Theory and Methods
@@
2.1 From micro- to macrostate dynamics
In a typical MSM workflow we first construct (say, $n\!\sim\!10^{2}$ $10^{3}$) microstates from the MD data, which are subsequently lumped into a few (say, $N\lesssim 10$) macrostates. As explained in the Introduction, we assume that the microstates show Markovian behavior at the chosen lag time $\tau$, which is sufficiently short to resolve the fastest dynamics of interest. Hence, the microstate transition matrix ${\boldsymbol{t}}=\{t_{ij}\}$ satisfies the Chapman-Kolmogorov relation31
@@ 2.1 From micro- to macrostate dynamics
This microstate-based evaluation of the macrostate transition matrix preserves all dynamical properties established for the microstates. For example, it yields by design exact macrostate population dynamics, and also provides correct state-to-state waiting times and transition pathways from a microstate Monte Carlo Markov chain simulations with subsequent projection on the macrostates. While the definition of ${\boldsymbol{T}}_{\rm Mic}$ is straightforward and its virtues are clearly promising, this microstate-based approach has to the best of our knowledge not yet been mentioned. Compared to the local-equilibrium approximation that yields a constant transition matrix ${\boldsymbol{T}}_{\rm LE}(\tau)$ and the MSM relation (8), however, the microstate-based evaluation results in a time-dependent macrostate transition matrix.
@@
Inserting the resulting ${\boldsymbol{K}}_{n}$ in Eq. (12), we obtain the desired evolution of ${\boldsymbol{T}}(t)$ for long times ($t\gg\tau_{\mathrm{K}}$). Hence the qMSM formulation facilitates a straightforward approximate solution of an in principle exact generalized master equation, and also explicitly provides the memory matrix ${\boldsymbol{K}}(t)$ that accounts for the non-Markovian dynamics of the system. Requiring only MD data of length $t\approx\tau_{\mathrm{K}}$, qMSM allows us to predict long-time dynamics from short trajectories. The above equations were implemented in an in-house Python code following Ref.  32; the kernel decay time $\tau_{\mathrm{K}}$ was obtained either directly from the memory kernel (in the one-dimensional toy model), or via the mean integral memory kernel17 (for HP35).
@@
Another issue is that qMSM keeps the additional term $\dot{{\boldsymbol{T}}}(0){\boldsymbol{T}}(t)$, which typically vanishes for symmetry reasons,16 but may be non-zero in a numerical evaluation. Interestingly, this term exactly cancels the memory kernel at $t\!=\!0$ [which can be shown by calculating ${\boldsymbol{T}}_{n}$ and ${\boldsymbol{K}}_{n}$ for $n=0$ from Eqs. (12) and (13), respectively]. While this is again a consequence of the chosen discretization, it seems odd, because ${\boldsymbol{K}}_{0}$ is the only non-zero term in the Markov limit. In all cases considered, however, the neglect of the term seems to hardly change the numerical results. Another practical issue is that the calculation of the memory matrix from Eq. (13) can be seriously plagued by noise resulting from the MD data and the time derivatives. As a remedy, recently an integrative ansatz 25 and a time-convolutionless approach 20 of the generalized master equation were proposed.
@@ 2.3 Hybrid MD/MSM formulation
In the Introduction, we opposed in Eqs. (1) and (2) two standard ways to estimate the macrostate transition matrix ${\boldsymbol{T}}(t)=\{T_{IJ}(t)\}$. On the one hand, we may use the MD data to directly count the transitions from state $J$ to state $I$ within time $t$. The resulting transition matrix ${\boldsymbol{T}}_{\rm MD}(t)$ generates the correct time evolution of the state model according to Eq. (1), but is limited to times $t\approx t_{\rm max}$ with $t_{\rm max}$ being the length of the MD trajectories (and actually even some hundreds MD frames less to facilitate a sufficient time average). On the other hand, we may invoke the Markov approximation to obtain ${\boldsymbol{T}}(t\!=\!m\tau)\approx{\boldsymbol{T}}^{m}(\tau)$, which assumes that the transition matrix ${\boldsymbol{T}}(t)$ becomes constant for $t\gtrsim\tau$. While this only requires short MD trajectories, we need to choose a lag time $\tau$ that is long enough to achieve Markovianity, but at the same time short enough to resolve the fastest dynamics of interest.
@@
In this way, we use at short times the approximation-free transition matrix ${\boldsymbol{T}}_{\rm MD}(t)$, including full time resolution (as given by the MD data). At long times, the MSM typically can use a rather long lag time.
@@ 3 Results
To compare the performance of the different methods, we test them on two model systems: A one-dimensional toy model lumping four microstates into two macrostates, and a recently established33 benchmark MSM of the folding of villin headpiece (aka HP35) using a long MD trajectory by Piana et al.30
@@ 3.1 One-dimensional toy model
${\boldsymbol{T}}_{\text{LE}}(t\!=\!m\tau_{0})=\begin{bmatrix}1-h/2&h/2\\ h/2&1-h/2\end{bmatrix}^{m}.$ (17)
@@
The Hummer-Szabo projection [Eq. (10)] gives ${\boldsymbol{T}}_{\rm HS}(t\!=\!m\tau_{0})={\boldsymbol{T}}_{\rm HS}^{m}$ with 13
@@
Figure 1: One-dimensional toy model. (a) Schematic free energy landscape, indicating four microstates 1, 2, 3 and 4, which are lumped into two macrostates $L$ and $R$. Depending on the ratio of the interstate transition probability h and the intrastate transition probability k, the macrostate dynamics is Markovian ($h/k\ll 1$, dark gray) or non-Markovian ($h/k\gtrsim 1$, dashed). (b) Ratio of the macroscopic and microscopic implied timescales $t_{\rm macro}/t_{\rm micro}$ of the two-macrostate model for $h/k=1$. Compared are the reference result obtained from the microstates (dashed black), the local equilibrium approximation (black), and the Hummer-Szabo projection (green), as well as the time evolution the result from qMSM (red) and of the microstate-based result (blue). (c) Normalized implied timescales obtained from the various methods, drawn as a function of the Markovianity parameter $h/k$. Time evolution of the macrostate population of the various methods, shown for (d) $h/k=1$ and (e) $h/k=5$. Also shown are results of the hybrid MD/MSM calculation [Eq. (14)] with $t_{\rm max}=10$. (f) Normalized memory kernel $K(t)$ of the qMSM method, obtained for $h/k=0.2$ (blue), $h/k=1$ (red), and $h/k=5$ (green).
As a first test, we wish to study how the various formulations reproduce the reference implied timescale $t_{\rm micro}$ of the microstates. Choosing $k=h=0.1$ (weakly non-Markovian case), Fig. 1b shows that the local equilibrium approximation $t_{\rm LE}$ underestimates $t_{\rm micro}$ almost by a factor 2, while the Hummer-Szabo result is only shorter by 13 %. The results obtained for ${\boldsymbol{T}}_{\text{Mic}}(t)$ and ${\boldsymbol{T}}_{\text{qMSM}}(t)$ reveal that the microstate-based method converges to the correct result and that qMSM converges to a value that is 9 % shorter than $t_{\rm micro}$. (In all cases, we choose the qMSM kernel decay time $\tau_{\rm K}$ such that the normalized memory kernel satisfies $K(\tau_{\rm K})\leq 0.1$.) To compare to a largely Markovian case ($h/k=0.2$), Fig. S1 shows that all methods yield the correct timescale $t_{\rm micro}$, except for the local equilibrium approximation, which gives $t_{\rm LE}=0.9\,t_{\rm micro}$.
We now choose $t=100$ (where qMSM and microstate-based results are virtually constant), and vary the Markovianity parameter $h/k$ from $h/k\ll 1$ (Markovian case) to $h/k\gtrsim 1$ (non-Markovian case). While for $h=0.1,\,k=0.5$ all methods yield the correct microstate timescale $t_{\rm micro}$ (except for $t_{\rm LE}$), the results vary considerably in the non-Markovian case. Choosing $h=0.5,\,k=0.1$, for example, $t_{\rm LE}$ underestimates $t_{\rm micro}$ by 87 %, $t_{\rm HS}$ by 40 %, and $t_{\rm qMSM}$ by 27 %, while the microstate-based result closely matches the reference.
@@
Choosing $h/k=1$ and $h/k=5$, Figs. 1d and 1e compare the population $P(t)$ obtained from the various formulations. While for $h/k=1$ all methods are quite close to the reference results, we find significant deviations in the strongly non-Markovian case $h/k=5$, where $P(t)$ in fact reveals two timescales. That is, the population exhibits a strong initial decay to $P(1)=0.75$ during the first time step, reflecting fast transitions over the shallow barrier at $x=0$ (Fig. 1a). All formulations reproduce this value correctly, except for the long-time approximation $P_{\rm HS}(t)$, which coincides with the reference for $t\gtrsim 7$. Being a short-time approximation, $P_{\rm LE}(t)$ decays significantly too fast for $t>1$. The qMSM calculation represents a clear improvement over these approximations and reproduces the exact results closely. Finally, we also show the hybrid calculation [Eq. (14)], which uses until $t_{\rm max}=10$ the exact transition matrix ${\boldsymbol{T}}_{\rm MD}(t)$, and for longer times the local approximation ${\boldsymbol{T}}_{\rm LE}(t=mt_{\rm max})$. For this choice of $t_{\rm max}$, the hybrid result virtually matches the reference at all times.